OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4).

OpenStudy (anonymous):

y=-4/3x

OpenStudy (anonymous):

first find slope of 4x-3y = -12 y = 4/3 x + 12 so slope is 4/3 and this is also the slope of the reqd line as they are parallel use general form: y-y1 = m(x-x1) where m = slope and a point on line is (x1,y1) so y-4 = (4/3) (x - (-3)) y - 4 = (4/3)x + 4 multiply through by 3 3y - 12 = 4x + 12 3y - 4x = 24 is the equation you want

OpenStudy (anonymous):

you can check this out by plugging in x = -3 , y = 4: 3(4) - 4(-3) = 12 + 12 = 24

OpenStudy (anonymous):

and if you rearrange the equation 3y = 4x + 24 y = (4/3) x + 8 - so the slope = 4/3